Grandma is 6 times as old as Erica. Grandma was 60 when Erica was born. How old is Grandma now?
Is this even possible to solve?

Answer by stanbon(75887)

Grandma is 6 times as old as Erica. Grandma was 60 when Erica was born. How old is Grandma now?
---------
Equations:
G = 6E
G = E + 60
------
Equation:
6E = E+60
5E = 60
E = 12 yrs (Erica's age now)
G = 6*E = 6*12 = 72 yrs (Grandma's age now)
====================
Cheers,
Stan H.
--------------------